# Harmonics of Learning: Universal Fourier Features Emerge in Invariant Networks

### v2

@@ 3 Theoretical Results
from which $\|W_{i}(1)\|=1$ follows. ∎
Again, the orthogonality assumption implies that $V_{1},\cdots,V_{k}$ are the ambient Hilbert spaces of all the irreducible unitary representations of $G$ up to isomorphism, and in particular $\sum_{i}\textnormal{dim}(V_{i})^{2}=|G|$.
Definition 8.
We say that $\varphi$ satisfies the adjunction property if
$\varphi(W,g\cdot x)=\varphi(g^{-1}\cdot W,x)$ (13)
for all $x\in\mathcal{X},g\in G$.
Remark 3.1.
As explained above, the tensor component $\langle G\rangle$ of $\mathcal{W}$ typically represents the input space of a given machine learning model. One can consider the more general scenario when data consist of complex signals over a finite set $\mathcal{S}$ acted upon by $G$, therefore replacing $\langle G\rangle=\mathbb{C}^{G}$ by $\mathbb{C}^{\mathcal{S}}$. This is the case, for example, for data consisting of images acted by the cyclic group via rotations, since the input space cannot be identified with signals over $G$. Assuming the action over $\mathcal{S}$ is free, meaning that $g\cdot s=s$ implies $g=1$, $\mathcal{S}$ can be decomposed into copies of $G$ deemed orbits. Specifically, there is an equivariant isomorphism $\mathcal{S}\simeq G\sqcup\cdots\sqcup G$, which in turn induces a linear isomorphism $\mathbb{C}^{\mathcal{S}}\simeq\langle G\rangle^{\oplus p}$, where $p$ is the number of orbits. The results from this section can be extended to this scenario by applying all the arguments to the copies of $\langle G\rangle$ separately, each of which will serve as a domain for its set of irreducible unitary representations. When the action is not free, it is necessary to take into account stabilizers, i.e. $g\in G$ such that $g\cdot s=s$ for some $s\in\mathcal{S}$. Roughly speaking, we expect that the results from this section can be adapted to an extent, obtaining unitary representations ‘up to stabilizers’. However, the precise meaning of the latter has yet to be clarified, and goes beyond the scope of this work.
3.1 Examples
In this section, we provide examples of machine learning models with unitary symmetries. As anticipated, in the context of machine learning $\mathcal{H}$ and $\mathcal{W}$ represent the hypothesis space and the parameter space, respectively. Indeed, in what follows $\mathcal{H}$ will consist of functions of the form $\langle G\rangle\rightarrow\mathcal{Y}$ for some codomain $\mathcal{Y}$, and we will adhere to the notation from Definition 8 accordingly. All the models considered in this section satisfy the adjunction property.
3.1.1 Spectral Networks
We start by considering Spectral Networks – a class of polynomial machine learning models that inspired this work—and to which our theory applies naturally. These models were introduced by [44] in cubic form based on the invariant theory of $\langle G\rangle$ (see Section B in the Appendix for an overview of the latter). The overall idea behind Spectral Networks is to approximate the $n$-order polynomial invariants (deemed spectral invariants) of $\langle G\rangle$ for an unknown group $G$. Specifically, suppose that $V_{1},\cdots,V_{k}$ are the ambient Hilbert spaces of the irreducible unitary representations of $G$. Given a multi-index $\underline{i}=(i_{1},\cdots,i_{n})\in\{1,\cdots,k\}^{n}$, The Spectral Network of order $n$ is defined as the collection of parametric maps $\varphi_{\underline{i}}(W,\cdot):\ \langle G\rangle\rightarrow\textnormal{End}(V_{i_{1}}\otimes\cdots\otimes V_{i_{n}})$:
where $W=\oplus_{i}W_{i}\in\mathcal{W}=\langle G\rangle\oplus_{i}\otimes\textnormal{End}(V_{i})$, $\odot$ denotes the $G$-wise tensor product of operators, and $\overline{x}$ denotes the component-wise conjugate of $x$. For a commutative $G$, since $V_{i}=\mathbb{C}$ for all $i$, the above expression reduces to:
$\varphi_{\underline{i}}(W,x)=W_{i_{1}}x\cdots\ W_{i_{n}}x\ \overline{W_{i_{1}}\odot\cdots\odot W_{i_{n}}\ x}.$ (15)
For $n=1$ Spectral Networks are deemed Power-Spectral Networks, and for a commutative $G$ they take the form $\varphi_{i}(W,x)=|W_{i}x|^{2}$. The latter can be simply interpreted as a linear model followed by an activation function. Even though the squared absolute value is uncommon as an activation function in machine learning, it has appeared in models of biological neural networks [1].
For simplicity, we will consider only the Spectral Networks involving a single unitary representation; that is, we will focus on constant multi-indices $\underline{i}=(i,\cdots,i)$ in Equation 14. To this end, let $V$ be a finite-dimensional Hilbert space and $\mathcal{H}$ be the set of functions $\langle G\rangle\rightarrow\textnormal{End}(V^{\otimes n})$ for some $n\in\mathbb{N}$. We set $\mathcal{W}=\langle G\rangle\otimes\textnormal{End}(V)$.
Proposition 3.3.
Consider the Spectral Network, given by:
$\varphi(W,x)=W(x)^{\otimes n}\ W^{\dagger\odot n}(\overline{x}).$ (16)
Then $\varphi$ has unitary symmetries.
We refer to the Appendix for a proof.
3.1.2 McCulloch-Pitts Neurons and Deep Networks
While Spectral Networks provide the most direct application of our theory, in this section we discuss the most common and fundamental neural network primitives in deep learning: the fully-connected McCulloch-Pitts neuron [31] and the deep neural network. We consider models with complex coefficients and focus on commutative groups, i.e. all the Hilbert spaces $V_{i}$ from Definition 7 will be equal to $\mathbb{C}$.
A McCulloch-Pitts neuron has the form $\varphi(W,x)=\sigma(Wx)$, where $\sigma\colon\mathbb{C}\rightarrow\mathcal{Y}$ is a map playing the role of an activation function and $W\in\mathcal{W}=\langle G\rangle$ is the weight vector. For $\sigma(z)=|z|^{2}$, the McCulloch-Pitts neuron reduces to a commutative Power-Spectral Network i.e., a Spectral Network with $n=1$. The hypothesis space $\mathcal{H}$ consists of functions $\langle G\rangle\rightarrow\mathcal{Y}$.
Proposition 3.4.
We refer to the Appendix for a proof. The above assumption on $\sigma$ is satisfied by popular activations functions from neuroscience and machine learning, such as the sigmoid and the leaky Rectified Linear Unit (ReLU), applied after taking complex absolute value.
Next, we discuss the case of classical deep networks. We model the latter as $\varphi(W,x)=\chi(|W(x)|^{2})$, where $W\in\mathcal{W}=\langle G\rangle\otimes\mathbb{C}^{k}$ and ||$\cdot$|| ⋅ | denotes the component-wise absolute value. Here, $k$ is the number of neurons in the first hidden layer of the network, while $\chi\colon\mathbb{R}_{\geq 0}^{k}\rightarrow\mathbb{R}$ is the head of the network, encompassing all the layers after the first one. Note that since typical neural networks are real-valued, we combine real and complex models. Namely, the first layer of $\varphi$ is complex, and its output is fed into the real head $\chi$ by taking squared absolute values.
Differently from Section 3, for the next result we will restrict $\mathcal{W}$ to the subspace of $\langle G\rangle\otimes\mathbb{C}^{k}$ consisting of $W$ such that the components $W_{i}$ are orthonormal. This implies, in particular, the constraint $k\leq|G|$. Note that $\mathcal{W}$ is closed by the actions of $G$ and $\textnormal{U}(\mathbb{C})$. The orthonormality condition is anyway necessary in order to recover the full Fourier transform, as stated in Corollary 3.2.
Proposition 3.5.
Suppose that there is an open subspace $O\subseteq\mathbb{R}_{\geq 0}^{k}$ containing 00 where $\chi$ is affine with distinct non-vanishing coefficients i.e., $\chi(z)=\sum_{i}a_{i}z_{i}+b$ for $z\in O$ with $0\not=a_{i}\not=a_{j}$ for $i\not=j$. Then $\varphi(W,x)=\chi(|W(x)|^{2})$ has unitary symmetries.
We refer to the Appendix for a proof. Since typical (real-valued) deep neural networks have piece-wise linear activations functions such as (leaky) ReLU, they define piece-wise affine maps and therefore are affine when restricted to appropriate open subspaces. Moreover, the hypothesis on the coefficients $a_{i}$ in Proposition 3.5 is generic, meaning that it defines an open dense subset of $(a_{1},\cdots,a_{k})\in\mathbb{R}^{k}$.
3.2 Group Recovery
Corollary 3.2 allows us to recover the group structure of $G$ up to isomorphism from the weights of a map with unitary symmetries. In other words, this enables the recovery of an unknown group in a data-driven manner from the weights of an invariant machine learning model, addressing the problem of symmetry discovery discussed in Section 1.1. The procedure was originally suggested and validated empirically in [44].
To this end, assume that $\varphi_{i}$ satisfies the requirements of Corollary 3.2. Moreover, we introduce the additional assumption that $W_{i}(1)=I\in\textnormal{End}(V_{i})$ for all $i$. If that is the case, $W$ coincides exactly with the Fourier transform by Corollary 3.2. This implies that the multiplication table of $G$ can be recovered by:
where $\odot$ denotes the Hadamard product i.e., component-wise operator composition. Note that this notation has a different meaning here than in Section 3.1.1. Since the $W(g)$’s are orthogonal, the only possible values for the norms over which the minimum is performed are 00 and $\sqrt{2|G|}$.
Definition 9.
$\omega_{\varphi}(\delta)=\sup_{\begin{subarray}{c}W,W^{\prime}\in\mathcal{W}\\ \|W\|=\|W^{\prime}\|\\ \Delta(\varphi(W),\varphi(W^{\prime}))\leq\delta\end{subarray}}\max_{i}\quad\inf_{U\in\textnormal{U}(V_{i})}\quad\|W_{i}-U\cdot W_{i}^{\prime}\|,$ (18)
Note that $\varphi$ has unitary symmetry if, and only if, $\omega_{\varphi}(0)=0$. The following is our main relaxation result.
Theorem 3.6.
Suppose that $\varphi\colon\mathcal{W}\rightarrow\mathcal{H}$ is a map and fix $W\in\mathcal{W}$. Denote
$L=\|W^{\dagger}W-|G|I\|_{\infty},$ (19)
where $\|\cdot\|_{\infty}$ is the uniform Frobenius norm for $G\times G$ matrices. Suppose that the following holds:
  • For all $g\in G$:
    $\omega_{\varphi_{i}}\left(\Delta(\varphi(g\cdot W),\varphi(W))\right)<\frac{\sqrt{\frac{1}{2}-\frac{L}{|G|}}}{\sqrt{|G|+L}+1}$ (20)
  • $L\leq\frac{|G|}{2}.$
  • $W_{i}(1)=I\in\textnormal{End}(V_{i})$ for all $i$.
Then Equation 17 holds, i.e. the group recovery algorithm is correct.
Proof.
@@ 3.2 Group Recovery
Firstly, the definition of $L$ implies the inequalities $|G|-L\leq\|W(g)\|^{2}\leq|G|+L$ and $|\langle W(g),W(h)\rangle|\leq L$ for all $g,h\in G$. In particular,
@@
By hypothesis, for each $i$ and $g\in G$ there exists $\rho_{i}(g)\in\textrm{U}(V_{i})$ such that
@@
Equation 23 implies that
@@
Since $\|W(p)-W(q)\|\geq\sqrt{2|G|-4L}$ for all $p\not=q\in G$, $W(g)\odot W(h)$ is closer to $W(gh)$ than to any other $W(q)$ for $q\in G$, which immediately implies the desired result. ∎
The assumption on $L$ in the above result is a relaxation of the unitarity assumption in Corollary 3.2 since $L=0$ if, and only if, $W$ is unitary up to a multiplicative constant.
Proposition 3.7.
Let $W\in\langle G\rangle$ and consider $\varphi(W,x)=\sigma(Wx)$. Suppose that the activation function $\sigma\colon\mathbb{C}\to\mathbb{C}$ is continuous and satisfies the following coercivity condition: there exist constants $C\in\mathbb{R}_{>0}$, $n\in\mathbb{N}$ such that for every $x\in\mathbb{C}$:
$|\sigma(0)-\sigma(x)|\geq C\ |x|^{n}.$ (28)
Then the unitarity defect of $\varphi$ satisfies for $\delta<C$:
3.2.1 Implementation
We now empirically explore the theory developed in this paper and demonstrate that Spectral Networks are able to recover the group structure in practice. To this end, we implement a non-commutative Power-Spectral Network $\varphi_{i}(W,x)=W_{i}x\ W^{\dagger}_{i}\overline{x}$ with weights $W\in\mathbb{C}^{d_{i}\times d_{i}\times d}$, where $d=|G|$ is the cardinality of the given group and $d_{1},\cdots,d_{k}$ are the dimensions of its irreducible unitary representations. As discussed at the beginning of Section 3.2, we force $W_{i}(1)=I\in\mathbb{C}^{d_{i}\times d_{i}}$, where the index $1$ is arbitrarily chosen.
Following [44], we train the model via contrastive learning [25]. Namely, given a finite dataset $\mathcal{D}$ of pairs $(x,y)$, where $x,y\in\langle G\rangle\simeq\mathbb{C}^{d}$ and $x=g\cdot y$ for an unknown $g\in G$, the objective optimized by the model is:
[…]
The model is trained via the Adam optimizer [28], which interprets the complex weights as real tensors of doubled dimensionality. The dataset $\mathcal{D}\subseteq\langle G\rangle\oplus\langle G\rangle$ is generated procedurally by first sampling $x$ from a standard Gaussian distribution over $\langle G\rangle\simeq\mathbb{R}^{2d}$, then sampling $g\in G$ uniformly, and finally producing the datapoint $(x,y=g\cdot x)\in\mathcal{D}$. We provide a Python implementation of both the model and the experiments at a public repository (see Section 1.1). The code is available in both the PyTorch [38] and the JAX [5] frameworks.
[…]
Figure 2: Learned Group Multiplication Tables. Tables inferred by Power-Spectral Networks for the groups $C_{6}$ (commutative), $C_{2}\times C_{2}\times C_{2}$ (commutative), and $D_{3}$ (non-commutative). Rows and columns are labeled with integers that index group elements. Each cell of the table contains the index of the group element obtained by composing the group elements indexed in the row and column. Note that the table is a symmetric matrix if, and only if, the group is commutative.
[…]
Figure 3: Group Recovery Accuracy. Accuracy for the recovery of the group multiplication tables across $20$ training runs as the amount of noise injected into data increases. The transition from $1.0$ to $0.0$ accuracy is sharp, and here we visualize only the noise regions where the values are non-trivial.
Once trained, we evaluate the model by checking whether the multiplication table $M\in\{1,\cdots,d\}^{d\times d}$ obtained via the group recovery algorithm described by Equation 17 coincides with the one of $G$. Since there is no canonical ordering on $G$, the table is recovered up to a permutation $\pi$ of $\{1,\cdots,d\}$ acting as $(\pi\cdot M)_{i,j}=\pi(M_{\pi^{-1}(i),\pi^{-1}(j)})$. Therefore, we check whether $\pi\cdot M$ coincides with the table of $G$ for all the permutations $\pi$. Figure 2 reports the (correct) multiplication tables obtained at convergence for both commutative and non-commutative groups. Specifically, we consider the cyclic group $C_{6}$, the product of cyclic groups $C_{2}\times C_{2}\times C_{2}$ and the dihedral group $D_{3}$, which is isomorphic to the symmetric group $S_{3}$.
In order to validate empirically the robustness of the group recovery procedure, we additionally train the model on data corrupted by white noise, i.e. $\mathcal{D}$ consists of pairs $(x,y=g\cdot x+\epsilon)$, where $\epsilon$ is sampled from an isotropic Gaussian distribution. We vary the standard deviation of the latter and report in Figure 3 the number of times the multiplication table is recovered correctly across $20$ training runs – a metric referred to as ‘table accuracy’.
As can be seen, the group structure is recovered most of the times even with large amounts of noise – up to $\sim 0.5$ standard deviation. The performance quickly degrades as the noise reaches a critical threshold. This demonstrates empirically that the group recovery procedure is robust to noise, which is in line with the theoretical bounds from Theorem 3.6.
4 Conclusions, Limitations, and Future Work
In this work, we proved that if a machine learning model of a certain kind is invariant to a finite group, then its weights are closely related to the Fourier transform on that group. We discussed how, as a consequence, the algebraic structure of an unknown group can be recovered from a model that is invariant. We established these results for both commutative and non-commutative groups, and discussed relaxed conditions under which the group recovery procedure holds. Our results represent a first step towards a mathematical explanation of universal features inferred by both biological and artificial neural networks.
In addition, our theoretical framework encompasses learning models with unitary symmetries. The latter is a technical property satisfied by Spectral Networks and, to an extent, by traditional deep networks. However, it is not clear what other models fit into our framework, and whether the notion is general enough to accommodate other computational primitives fundamental to machine learning, such as the attention mechanisms. This is an open question that is worthy of investigation.
Acknowledgements
Appendix
Appendix A Proofs of Theoretical Results
A.1 Proof of Proposition 3.3
Lemma A.1.
Proof.
From the polar decomposition for matrices (see [18, Theorem 3.1.9]), we know that:
$A=PV,\ B=QW,$ (31)
where $P,Q\in\mathbb{C}^{d^{\prime}\times d^{\prime}}$ are Hermitian positive semidefinite and $V,W\in\mathbb{C}^{d^{\prime}\times d}$ have orthonormal rows. Also, $P^{2}=AA^{\dagger}=BB^{\dagger}=Q^{2}$ from which it follows that $P=Q$ by uniqueness of square roots of positive semidefinite Hermitian matrices (see [19, Theorem 7.2.6]). In particular, we have:
$A=PV=QV=QWW^{\dagger}V=BU,$ (32)
with $U=W^{\dagger}V$ unitary. ∎
Remark A.1.
Lemma A.2.
Proof.
From the Spectral Theorem we know that:
$A=UDU^{\dagger},$ (33)
where $U$ is unitary and $D$ is diagonal. From $A^{\otimes n}=B^{\otimes n}$ it follows that $D^{\otimes n}=C^{\otimes n}$, where $C=U^{\dagger}BU$. Note that the (point-wise) Hadamard product of matrices is a submatrix of the tensor (Kronecker) product. In particular, the off-diagonal entries of $C$ must vanish. On the diagonal we have $(D_{i,i})^{n}=(C_{i,i})^{n}$ for every $i$, and therefore $D_{i,i}=C_{i,i}$ since they are non-negative. This shows that $C=D$, implying that $B=UDU^{\dagger}=A$. ∎
By putting together the above lemmas, we obtain the following.
Lemma A.3.
Let $A_{1},\cdots,A_{k},B_{1},\cdots,B_{k}\in\mathbb{C}^{d\times d}$. Suppose that for some $n>0$, for all $x\in\mathbb{C}^{k}$:
Proof.
$\sum_{i_{1},\cdots,i_{n+1}}x_{i_{1}}\cdots x_{i_{n}}\overline{x}_{i_{n+1}}(A_{i_{1}}\otimes\cdots\otimes A_{i_{n}})A_{i_{n+1}}^{\dagger\otimes n}=$ (35)
$=$ $\sum_{i_{1},\cdots,i_{n+1}}x_{i_{1}}\cdots x_{i_{n}}\overline{x}_{i_{n+1}}(A_{i_{1}}A_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(A_{i_{n}}A_{i_{n+1}}^{\dagger})=$ (36)
$=$ $\sum_{i_{1},\cdots,i_{n+1}}x_{i_{1}}\cdots x_{1_{n}}\overline{x}_{i_{n+1}}(B_{i_{1}}B_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(B_{i_{n}}B_{i_{n+1}}^{\dagger}).$ (37)
Since a polynomial vanishes as function if and only if it is the zero polynomial, it follows that $(A_{i_{1}}A_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(A_{i_{n}}A_{i_{n+1}}^{\dagger})=(B_{i_{1}}B_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(B_{i_{n}}B_{i_{n+1}}^{\dagger})$ for all $i_{1},\cdots,i_{n+1}$, and in particular $(A_{i}A_{j}^{\dagger})^{\otimes n}=(B_{i}B_{j}^{\dagger})^{\otimes n}$ for all $i,j$. From Lemma A.2 we conclude that $A_{i}A_{j}=B_{i}B_{j}$ for all $i,j$, which can be rephrased as $AA^{\dagger}=BB^{\dagger}$, where $A,B$ are the $(dk)\times d$ matrices obtained by concatenating the $A_{i}$’s and $B_{i}$’s respectively. From Lemma A.1 we conclude that $A=BU$ for a unitary $d\times d$ matrix $U$, as desired. ∎
A.2 Proof of Proposition 3.4
Proof.
Pick $W,W^{\prime}\in\langle G\rangle$ of the same norm such that $\varphi(W,x)=\varphi(W^{\prime},x)$ for all $x\in\langle G\rangle$. Given the open set $O\subseteq\mathbb{C}$ from the hypothesis on $\sigma$, consider $O^{\prime}=\{x\in\langle G\rangle\ |\ Wx,W^{\prime}x\in O\}$, which is open and non-empty since $0\in O^{\prime}$. For $x\in O^{\prime}$, $Wx=0$ implies $\varphi(W,x)=\varphi(W^{\prime},x)=0$, from which it follows that $W^{\prime}x=0$ by definition of $O$. Therefore, $W$ and $W^{\prime}$ share the same orthogonal complement, implying that $W^{\prime}=\rho W$ for some $\rho\in\mathbb{C}$. Since $W$ and $W^{\prime}$ have the same norm, we conclude that $\rho\in\textnormal{U}(\mathbb{C})$. ∎
A.3 Proof of Proposition 3.5
Proof.
Consider $W,W^{\prime}\in\mathcal{W}$ such that $\varphi(W,x)=\varphi(W^{\prime},x)$ for all $x\in\langle G\rangle$. Given the open set $O\subseteq\mathbb{R}_{\geq 0}^{k}$ from the hypothesis on $\chi$, consider $O^{\prime}=\{x\in\langle G\rangle\ |\ |Wx|^{2},|W^{\prime}x|^{2}\in O\}$, which is open and non-empty since $0\in O^{\prime}$. For $x\in O^{\prime}$, $\varphi(W,x)$ can be written as:
$\varphi(W,x)=\sum_{i}a_{i}|W_{i}\ x|^{2}+b,$ (38)
with the $a_{i}$’s being distinct, and similarly for $\varphi(W^{\prime},x)$. Since Hermitian forms are determined by their restriction on an open set, we deduce the following identity of operators:
$\sum_{i}a_{i}W_{i}\otimes\overline{W_{i}}=\sum_{i}a_{i}W_{i}^{\prime}\otimes\overline{W_{i}^{\prime}}.$ (39)
Since both the sets $\{W_{i}\}_{i}$ and $\{W_{i}^{\prime}\}_{i}$ are orthonormal by hypothesis on $\mathcal{W}$, both sides of Equation 39 define a spectral decomposition, i.e. a decomposition into projections over orthonormal vectors. Since the eigenvalues $a_{i}$ are distinct and non-vanishing, it follows that $W_{i}=\rho_{i}W_{i}^{\prime}$ for some $\rho_{i}\in\textnormal{U}(\mathbb{C})$, as desired. ∎
A.4 Proof of Proposition 3.7
@@ Appendix A Proofs of Theoretical Results
Let $H$ be a finite-dimensional complex Hilbert space, $v,w\in H$ normal and $\varepsilon\in\mathbb{R}$ such that $10<\varepsilon<$ 1. Suppose that for every normal $x$ orthogonal to $w$, it holds that $|\langle x,v\rangle|\leq\varepsilon$. Then there exists $\rho\in\textnormal{U}(\mathbb{C})$ such that
@@
The hypothesis implies then that $\ \langle w_{2},v\rangle\ \geq\sqrt{1-\varepsilon^{2}}$. Therefore,
@@
By Lemma A.4 there exists $\rho\in\textrm{U}(\mathbb{C})$ such that:
@@
from which the claim follows. ∎

### v3

@@ Appendix A Proofs of Theoretical Results
from which $\|W_{i}(1)\|=1$ follows.
A.3 Proof of Proposition 3.6
Lemma A.3.
Proof A.4.
From the polar decomposition for matrices (Horn et al., 1994, Theorem 3.1.9), we know that:
$A=PV,\ B=QW,$ (15)
where $P,Q\in\mathbb{C}^{d^{\prime}\times d^{\prime}}$ are Hermitian positive semidefinite and $V,W\in\mathbb{C}^{d^{\prime}\times d}$ have orthonormal rows. Also, $P^{2}=AA^{\dagger}=BB^{\dagger}=Q^{2}$ from which it follows that $P=Q$ by uniqueness of square roots of positive semidefinite Hermitian matrices (Horn and Johnson, 2012, Theorem 7.2.6). In particular, we have:
$A=PV=QV=QWW^{\dagger}V=BU,$ (16)
with $U=W^{\dagger}V$ unitary.
Remark A.5.
Lemma A.6.
Proof A.7.
From the Spectral Theorem we know that:
$A=UDU^{\dagger},$ (17)
where $U$ is unitary and $D$ is diagonal. Since $A^{\otimes n}=B^{\otimes n}$, it follows that $D^{\otimes n}=C^{\otimes n}$, where $C=U^{\dagger}BU$. Note that the (point-wise) Hadamard product of matrices is a submatrix of the tensor (Kronecker) product. In particular, the off-diagonal entries of $C$ must vanish. On the diagonal we have $(D_{i,i})^{n}=(C_{i,i})^{n}$ for every $i$, and therefore $D_{i,i}=C_{i,i}$ since they are non-negative. This shows that $C=D$, implying that $B=UDU^{\dagger}=A$.
Putting together the above lemmas, we obtain the following.
Lemma A.8.
Let $A_{1},\ldots,A_{k},B_{1},\ldots,B_{k}\in\mathbb{C}^{d\times d}$. Suppose that for some $n>0$, for all $x\in\mathbb{C}^{k}$:
Proof A.9.
$\sum_{i_{1},\ldots,i_{n+1}}x_{i_{1}}\cdots x_{i_{n}}\overline{x}_{i_{n+1}}(A_{i_{1}}\otimes\cdots\otimes A_{i_{n}})A_{i_{n+1}}^{\dagger\otimes n}=$ (19)
$=$ $\sum_{i_{1},\ldots,i_{n+1}}x_{i_{1}}\cdots x_{i_{n}}\overline{x}_{i_{n+1}}(A_{i_{1}}A_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(A_{i_{n}}A_{i_{n+1}}^{\dagger})=$ (20)
$=$ $\sum_{i_{1},\ldots,i_{n+1}}x_{i_{1}}\cdots x_{1_{n}}\overline{x}_{i_{n+1}}(B_{i_{1}}B_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(B_{i_{n}}B_{i_{n+1}}^{\dagger}).$ (21)
Since a polynomial vanishes as function if and only if it is the zero polynomial, it follows that $(A_{i_{1}}A_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(A_{i_{n}}A_{i_{n+1}}^{\dagger})=(B_{i_{1}}B_{i_{n+1}}^{\dagger})\otimes\cdots\otimes(B_{i_{n}}B_{i_{n+1}}^{\dagger})$ for all $i_{1},\ldots,i_{n+1}$, and in particular $(A_{i}A_{j}^{\dagger})^{\otimes n}=(B_{i}B_{j}^{\dagger})^{\otimes n}$ for all $i,j$. From Lemma A.6 we conclude that $A_{i}A_{j}=B_{i}B_{j}$ for all $i,j$, which can be rephrased as $AA^{\dagger}=BB^{\dagger}$, where $A,B$ are the $(dk)\times d$ matrices obtained by concatenating the $A_{i}$’s and $B_{i}$’s, respectively. From Lemma A.3 we conclude that $A=BU$ for a unitary $d\times d$ matrix $U$, as desired.
A.4 Proof of Proposition 3.7
Proof A.10.
Pick $W,W^{\prime}\in\langle G\rangle$ of the same norm such that $\varphi(W,x)=\varphi(W^{\prime},x)$ for all $x\in\langle G\rangle$. Given the open set $O\subseteq\mathbb{C}$ from the hypothesis on $\sigma$, consider $O^{\prime}=\{x\in\langle G\rangle\ |\ Wx,W^{\prime}x\in O\}$, which is open and non-empty since $0\in O^{\prime}$. For $x\in O^{\prime}$, $Wx=0$ implies $\varphi(W,x)=\varphi(W^{\prime},x)=0$, from which it follows that $W^{\prime}x=0$ by definition of $O$. Therefore, $W$ and $W^{\prime}$ share the same orthogonal complement, implying that $W^{\prime}=\rho W$ for some $\rho\in\mathbb{C}$. Since $W$ and $W^{\prime}$ have the same norm, we conclude that $\rho\in\textnormal{U}(\mathbb{C})$.
A.5 Proof of Proposition 3.9
Proof A.11.
Pick $W,W^{\prime}\in\mathcal{W}$ such that $\|W_{i}\|=\|W_{i}^{\prime}\|$ for all $i$ and $\varphi(W,x)=\varphi(W^{\prime},x)$ for all $x\in\langle G\rangle$. Given the open set $O\subseteq\mathbb{R}^{k}$ from the hypothesis on $\chi$, consider the set $O^{\prime}=\{x\in\langle G\rangle\ |\ \sigma(W(x)),\sigma(W^{\prime}(x))\in O\}$, which is open and non-empty since $0\in O^{\prime}$. For $x\in O^{\prime}$, an immediate calculation shows that:
$\partial_{x}\partial_{\overline{x}}\ \varphi(W,x)=\sum_{i}a_{i}\ \partial_{z}\partial_{\overline{z}}\sigma(W_{i}\ x)\ W_{i}\otimes\overline{W_{i}},$ (22)
with the $a_{i}$’s being distinct, and similarly for $\varphi(W^{\prime},x)$. By specializing to $x=0$, we deduce that:
[…]
Since both the sets $\{W_{i}\}_{i}$ and $\{W_{i}^{\prime}\}_{i}$ are orthonormal by hypothesis on $\mathcal{W}$, both sides of Equation 23 define a spectral decomposition, i.e. a decomposition into projections over orthonormal vectors. From the hypothesis on $\chi$ it follows that the eigenvalues $a_{i}\ \partial_{z}\partial_{\overline{z}}\sigma(W_{i}\ x)$ of the operators in Equation 23 are distinct and non-vanishing, implying that the eigenvectors coincide up to a multiplicative scalar, i.e. $W_{i}=\rho_{i}W_{i}^{\prime}$ for some $\rho_{i}\in\textnormal{U}(\mathbb{C})$, as desired.
A.6 Proof of Theorem 3.11
Proof A.12.
@@
Firstly, the definition of $L$ implies the inequalities $|G|-L\leq\|W(g)\|^{2}\leq|G|+L$ and $|\langle W(g),W(h)\rangle|\leq L$ for all $g,h\in G$. In particular, we have:
@@
By hypothesis, for every $i$ and $g\in G$ there exists $\rho_{i}(g)\in\textrm{U}(V_{i})$ such that:
@@
Equation 26 implies that:
@@
Since $\|W(p)-W(q)\|\geq\sqrt{2|G|-4L}$ for all $p\not=q\in G$, $W(g)\odot W(h)$ is closer to $W(gh)$ than to any other $W(q)$ for $q\in G$, which immediately implies the desired result.
A.7 Proof of Proposition 3.12
@@
Let $H$ be a finite-dimensional complex Hilbert space, $v,w\in H$ normal and $\varepsilon\in\mathbb{R}$ such that $10<\varepsilon<$ 1. Suppose that for every normal $x$ orthogonal to $w$, it holds that $|\langle x,v\rangle|\leq\varepsilon$. Then there exists $\rho\in\textnormal{U}(\mathbb{C})$ such that:
@@
The hypothesis implies then that $\ \langle w_{2},v\rangle\ \geq\sqrt{1-\varepsilon^{2}}$. Therefore, we have:
@@
By Lemma A.13, there exists $\rho\in\textrm{U}(\mathbb{C})$ such that:
@@
from which the claim follows.